To: J. Cook, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IR Demo Project Weekly Report, April 20 - 24, 1998
Date: April 24, 1998
This week involved a full week of commissioning
activities with a scheduled maintenace shift on Tuesday, April
21. Highlights for the week included some additional running
at mA levels to look at cathode lifetime issues, and continued
check-out of electron beam diagnostics to quantify the beam quality.
We got our first look at some of the beam parameters after the
injector (emittance, bunch length, energy spread). The first values
are within factors of 2-3 of the desired values, which is not
bad for our first measurements in a non-optimized system. (See
Commissioning Report below for details.)
The March 1998 Monthly Reports for the project
were distributed to the DOE and Navy Program Offices on Wednesday
On Tuesday, April 21, the new acting Head
of DOE's Office of Nuclear Physics, Dr. Dennis Kovar, and Sherman
Fivozinski, the Jefferson Lab Contract Monitor for DOE-NP visited
the Lab for a general briefing that included a tour of the FEL
The MPS was updated this week to include the magnet to send the
beam to the 0G dump or the cryomodule. The 2F06 and 2G00 BLMs
were mapped. The 0G01 BLM is functional, but not yet mapped.
The 180° Bend Magnet (DY) is in the final set of measurements.
All earlier measurements looked fine. The first article DG corrector
has been received and measured to be within its specifications.
A possible solution for the problem with the Drive Laser Pulse
Controller will be tested this weekend.
The first FEL Mirror Cassette is at assembly. The FEL pipe is
aligned in the Laser Control room and fiducials are established
for the User Labs. Control cabling for the Laser Control room
is in progress.
The ND2 filter for the Drive Laser was placed under EPICS control.
The installation of a CARM for the Wiggler area is in progress.
This week was devoted principally to measurements of beam quality,
as well as improving the injector setup, but it also included
a few cw runs up to mA-level average current at 38 MeV energy
to the straight-ahead dump. We are generally measuring six-dimensional
phase-space parameters of the beam that are within factors of
2-3 of theoretical, a favorable situation given that we just turned
on the machine. It means there are no apparent show-stoppers
to lasing, but fine tuning is needed beforehand. Specific measurements
performed this week include rms normalized transverse emittance
at the cryounit exit (7-8 mm-mrad vs. 3.5 mm-mrad from PARMELA),
rms bunch length at the center of the cryomodule (about 2 ps,
vs. 1 ps desired at the wiggler), rms energy spread at the center
of the first optical chicane (0.5% vs. 0.15% desired at the wiggler),
and a rough sense that the emittance at the wiggler is not too
far from the 9 mm-mrad desired in that most of the electron beam
can be threaded through the mm- size holes in the three wiggler
viewers. Getting a sufficient array of beam diagnostics working
and using them to configure the quality of the electron beam at
the wiggler location are now our highest-priority tasks. They
will occupy the bulk of our commissioning time in the next two
weeks, in which will be interspersed some cw runs at high average
current to exercise the machine and get more data on cathode lifetime.